OpenStudy (anonymous):

A computer is purhcased for 980 pounds.Its value will decrease by 4% per year.How much will it be worth in 2 years time.

OpenStudy (anonymous):

Can you tell me its value at the end of 1 year?

OpenStudy (anonymous):

thats what i am struggling on @Mashy

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

after the first year the value of the computer is: \[\Large 980 - \left( {\frac{4}{{100}} \cdot 980} \right) =940.8\] and after the second year the value of the computer is: \[\Large 940.8 - \left( {\frac{4}{{100}} \cdot 940.8} \right) = ...?\]

OpenStudy (anonymous):

the second year is 37.632 i think , I did 4/100 multiplied by 940.80

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

am I also meant to subtract somewhere @Michele_Laino

OpenStudy (michele_laino):

now, you have to do this subtraction: \[\Large 970.8 - 37.63 = ...?\]

OpenStudy (michele_laino):

oops.. \[\Large 940.8 - 37.63 = ...?\]

OpenStudy (anonymous):

903.17

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

is that the answer?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

how did you get 4/100 multiplied by 980 as 940.8 @Michele_Laino

OpenStudy (anonymous):

i got 39.2 @Michele_Laino

OpenStudy (michele_laino):

since we have to do this subtraction: \[\Large 980 - \left( {\frac{4}{{100}} \cdot 980} \right) = 980 - 39.2 = 940.8\]

OpenStudy (anonymous):

why do we have to do a subtraction ?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

because the value of the computer is decreasing from its original value

OpenStudy (anonymous):

why did we have to do 4 percent of 940.8 as well as 4 percent of 980 @Michele_Laino

OpenStudy (anonymous):

thanks for the reply for my first query

OpenStudy (michele_laino):

since at the end of one year, the value of the computer is decreased by 4% from its starting value. Now the starting value at the first year is $980.00, whereas the starting value at the beginning of the second year is $940.80

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

:)